KyledaGreat:

Consider the following function. Graph the original function by indicating how the more basic function has been shifted, reflected, stretched, or compressed.

KyledaGreat:

1 attachment
KyledaGreat:

1 attachment
KyledaGreat:

1 attachment
KyledaGreat:

@vocaloid @surjithayer @snowflake0531

KyledaGreat:

@vocaloid @surjithayer @snowflake0531

surjithayer:

\[p(x)=(x-4)^3\] is this the original function?

surjithayer:

or original function is \[f(x)=x^3\]

surjithayer:

sorry \[p(x)=x^3\]

surjithayer:

if original function is \[p(x)=x^3~~~~...(1)\] new functin is \[p(x)=(x-4)^3~~~~...(2)\] (2) is obtained by 4 units Horizontal shift of (1)

KyledaGreat:

the horizontal shift is left ?

surjithayer:

if (1) is original function ,then (2) is Horizontal shift by 4 units right.

surjithayer:

vrtical stretch-none

surjithayer:

no reflection on x-axis or y-axis.

KyledaGreat:

okay , the horizontal shift is right and 4 units?

KyledaGreat:

and vertical shirt is ?

surjithayer:

none

KyledaGreat:

shift sorry and okay

surjithayer:

if function is \[p(x)=(x-4)^3+p\] if p is positive,then vertical shift up if p is negative then vertical shift is down. if p=0,then no vertical shift.

KyledaGreat:

it's up

KyledaGreat:

am i right ?

KyledaGreat:

am i right ?

surjithayer:

function is \[p(x)=(x-4)^3\] it has no vertical shift.

KyledaGreat:

it was right and 4 , it was wrong

surjithayer:

what you mean?

KyledaGreat:

everything was right but i put in something else

surjithayer:

am i correct?